# FedSynthCT-Brain: A Federated Learning Framework for Multi-Institutional Brain MRI-to-CT Synthesis

### v1

@@ 5.2 Benchmarking aggregation strategies
Table 4 presents the best result obtained for each aggregation strategy, tested on the unseen dataset (centre E). Results are reported in terms of MAE, which resulted the most sensitive image similarity metric in the benchmarking of DL architectures. The experiments (Experiment 1, 2, 3, 4, 5), were repeated 5 times for each strategy to determine the range of variability of the convergence trend within 25 rounds. This comparison helps to understand which aggregation strategy gives the best result in a stable and efficient way. For each experiment, the number of rounds performed to obtain the best result is also reported.
Table 4: Results obtained for the tested aggregation strategies in terms of image similarity and the number of rounds required to achieve the best result. The mean and standard deviation (Std) are presented for both rounds and the Mean Absolute Error (MAE).
@@ 5.3 Federated sCTs
Figure 6: The figure shows the centre slice for the axial, coronal and sagittal planes, which is used for an evaluation of the worst and best –A and B respectively– predictions based on image similarity metrics of the federated model on the unseen dataset (Centre E). In the case B, the red square highlights a calcification in the ground truth CT, which is not correctly predicted by the model. In the case B, the blue rectangle also highlights a thin duplication of the upper skull bone. This issue has been identified exclusively within the predictions generated by the federated model on the Centre E dataset. The colour map on the left illustrates the differences between the CT and the sCT, highlighting regions where the model exhibits a higher or lower tendency to over or underestimate the actual outcome.
Figure 7: The figure shows the central slice for the axial, coronal, and sagittal planes, which is used for an assessment of one of the two test cases of Centres A and B, respectively. The red boxes in the Centre A-coronal view-comparison highlight the approximation of the small structure performed by the model during the sCT generation. The blue boxes in the Centre B-coronal and sagittal views-comparison indicate the presence of an artefact in the ground truth CT, whereby the upper skull is larger than expected. Despite this artefact, the federated model predicts a smaller skull structure, and thus a more realistic outcome.
@@ 6 Discussion
Regarding the evaluation of U-Net architectures, the results obtained on Centre E — the unseen dataset — as well as across the various Centres in the federation, reveal no relevant differences in image similarity metrics when comparing different architectures. Based on the MAE metric, the minimal variation in HU does not allow for the identification of a clearly superior U-Net design. This could imply that, within the context of the FedSynthCT-Brain framework, the architectures are sufficiently robust to handle the data variability across the different centres. Consequently, the focus shifted to evaluating the disparity in efficiency among the different implementations, as this becomes a critical factor in determining the most suitable model for FL applications.
@@
We also integrated the winning MSEP architecture from the SynthRAD2023 challenge into our FL framework by adapting its original 3D implementation. Despite efforts to reduce computational overhead — such as limiting the number of participating clients and reducing the size of the 3D patches — the time required for a single federated round remained prohibitively high. Specifically, one round took approximately 85 hours and 37 minutes, compared to just 49 minutes for the U-Net model proposed by Li et al. [34]. Moreover, the federated MSEP architecture’s performance fell short when compared to its centralised counterpart. This decline can be primarily attributed to: (i) the adjustments made to the training pipeline to accommodate the federated framework, and (ii) the reduced number of rounds and clients necessitated by the computational complexity of the 3D approach.
Moving to the aggregation strategies, the analysis of FedAvg results (see Table 4, revealed that the optimal performance were obtained in 25 rounds, exhibiting minimal variation – $1.4$ HU – in terms of MAE. However, when compared to the proposed aggregation method (FedAvg+FedProx), FedAvg required an average higher number of rounds, indicating a longer time to reach its optimal result. Referring to the Table 4, using FedProx enhances the performance of FedAvg and accelerates the convergence process, achieving its optimal result in an average of $12.6\pm 3.7$ rounds, compared to $14.8\pm 4.1$ rounds for the FedAvg alone. This finding aligns with previous studies in closer fields of research (e.g., [33]), which highlight that the introduction of a proximal term can mitigate the impact of client heterogeneity and improve convergence stability in FL settings.
@@
Some inconsistencies were identified in the upper skull of a few patients, indicated by the blue box in Figure 6, case B). The wrong I2I translation produced a subtle bone mirroring. In other cases the federated model was not able to predict the presence of calcifications in the brain tissue. This failure can be attributed to the limited number of examples that exhibited this brain calcification in the federation, which precluded the model from identifying abnormal cases such as the one highlighted.
Despite the Centre E’s dataset being external to the training, the federated model demonstrated its capacity to predict a missing portion of bone in the axial view (see Figure 6, case A, green box).
@@
As described in Section 3.1, Centre B was introduced in the federation to increase data heterogeneity, as the MRIs were acquired with a hybrid PET-MRI scanner, with larger CT slice thickness and different X-ray tube energy (120 kV vs. 140 kV). From Table 2, it is evident that the error for the Centre B is greater than for the other clients, emphasising the difference from the other images in the federation. However, it can be observed that the error is partly due to the lower spatial resolution of the ground truth CT, introducing partial volume artefacts in the upper part of the skull (see Figure 7, light blue box). The result of this artefact is a noticeable, unrealistict thickening of the bone, as presented in Figure 9. Despite this problem, the skull bone of the predicted federated sCT appears more realistic and consistent with other examples. This establishes an advantage of the federation, as the influence given by the weights of the other Centres allows the federated model to remain robust in front of optimization errors during training.
@@
This work established the foundations for future developments. A possible improvement could involve optimizing the implementation of 3D architectures for federated scenarios by leveraging more efficient training strategies, such as model pruning, knowledge distillation, or dimensionality reduction [48]. Additionally, exploring hybrid federated approaches, such as cross-silo horizontal FL with asynchronous updates or client clustering [40, 57], could reduce synchronization times.
Moreover, the implementation of specific pre-processing or data harmonisation techniques can be considered to enhance the homogeneity of the data and to mitigate potential hidden artefacts. This, consequently, will improve the quality of the sCTs for both the Centres within the federation and in the generalisation task.
@@ Acknowledgments
Appendix A Centre B artefact measurement
@@
Appendix B Aggregation strategies convergence trends

### v2

@@ 5.2 Benchmarking aggregation strategies
Table 3 presents the best result obtained for each aggregation strategy, tested on the unseen dataset (centre E). Results are reported in terms of MAE, which resulted the most sensitive image similarity metric in the benchmarking of DL architectures. The experiments (Experiment 1, 2, 3, 4, 5), were repeated 5 times for each strategy to determine the range of variability of the convergence trend within 25 rounds. This comparison helps to understand which aggregation strategy gives the best result in a stable and efficient way. For each experiment, the number of rounds performed to obtain the best result is also reported. Figure 10 reports the convergence plots for all evaluated strategies. Furthermore, a visual comparison of the results obtained with different strategies were reported in Figure 11.
Table 3: Comparison, in terms of image similarity (MAE) and number of rounds to achieve the best results among different aggregation strategies. Experiments were repeated 5 times and the mean and standard deviation (Std) are presented for both number of rounds and the MAE.
@@ 5.3 Federated sCTs
Figure 6: The figure shows the central slice for the axial, coronal and sagittal planes, which is used for an evaluation of the worst and best –A and B respectively– predictions based on image similarity metrics of the federated model on the unseen dataset (Centre E). The Detail A (in green), shows the ability of the model to predict a missing portion of bone, although the case A was from the unseen Centre. In the case B, the Detail B.1 (in red) highlights a calcification in the ground truth CT, which is not correctly predicted by the model. In the case B, the Detail B.2 (in blue) also highlights a thin duplication of the upper skull bone. This issue has been identified exclusively within the predictions generated by the federated model on the Centre E dataset. The colour map on the left illustrates the differences between the CT and the sCT, highlighting regions where the model exhibits a higher or lower tendency to over or underestimate the actual outcome.
Figure 7: The figure shows the central slice for the axial, coronal, and sagittal planes, which is used for an assessment of one of the two test cases of Centres A and B, respectively. The Detail A (in red) comparison highlight the approximation of the small structure performed by the model during the sCT generation for Centre A. Detail B comparison (in blue) shows the presence of artefact in Centre B’s ground-truth CTs, whereby the upper skull is larger than expected. Despite this artefact, the federated model predicts a smaller skull structure, and thus a more realistic outcome.
@@ 6 Discussion
Regarding the evaluation of U-Net architectures, the results obtained on Centre E — the unseen dataset — as well as across the various Centres in the federation, reveal no relevant differences in image similarity metrics when comparing different architectures. Based on the MAE metric, the minimal variation in HU does not allow for the identification of a clearly superior U-Net design. This could imply that, within the context of the FedSynthCT-Brain framework, the architectures are sufficiently robust to handle the data variability across the different centres. Furthermore, SSIM and PSNR, being less sensitive to small variations than MAE, remained stable when comparing different architectures within the same centre. Consequently, the focus shifted to evaluating the disparity in efficiency among the different implementations, as this becomes a critical factor in determining the most suitable model for FL applications.
@@
The employment of U-Net architectures was further justified by supplementary experiments comparing various models, including Conditional GAN, CycleGAN and Transformers. The results of these experiments, which are discussed in A, demonstrated that the U-Net achieved the optimal overall performance, balancing image similarity metrics and computational efficiency, thus rendering them more suitable for FL MRI-to-sCT applications.
Moving to the aggregation strategies, the analysis of FedAvg results (see Table 3, revealed that the optimal performance were obtained in 25 rounds, exhibiting minimal variation – $1.4$ HU – in terms of MAE. However, when compared to the proposed aggregation method (FedAvg+FedProx), FedAvg required an average higher number of rounds, indicating a longer time to reach its optimal result. Referring to the Table 3, using FedProx enhances the performance of FedAvg and accelerates the convergence process, achieving its optimal result in an average of $12.6\pm 3.7$ rounds, compared to $14.8\pm 4.1$ rounds for the FedAvg alone. This finding aligns with previous studies in closer fields of research, which highlight that the introduction of a proximal term can mitigate the impact of client heterogeneity and improve convergence stability in FL settings.
@@
Some inconsistencies were identified in the upper skull of a few patients, indicated by the Detail B.2 in Figure 6, case B). The wrong I2I translation produced a subtle bone mirroring. In other cases the federated model was not able to predict the presence of calcifications in the brain tissue. This failure can be attributed to the limited number of examples that exhibited this brain calcification in the federation, which precluded the model from identifying abnormal cases such as the one highlighted.
Despite the Centre E’s dataset being external to the training, the federated model demonstrated its capacity to predict a missing portion of bone in the axial view (see Figure 6, case A, Detail A).
@@
As described in Section 3.1, Centre B was introduced in the federation to increase data heterogeneity, as the MRIs were acquired with a hybrid PET-MRI scanner, with larger CT slice thickness and different X-ray tube energy (120 kV vs. 140 kV). From Table 2, it is evident that the error for the Centre B is greater than for the other clients, emphasising the difference from the other images in the federation. However, it can be observed that the error is partly due to the lower spatial resolution of the ground truth CT, introducing partial volume artefacts in the upper part of the skull (see Figure 7, Detail B). The result of this artefact is a noticeable, unrealistict thickening of the bone, as presented in Figure 9. Despite this problem, the skull bone of the predicted federated sCT appears more realistic and consistent with other examples. This establishes an advantage of the federation, as the influence given by the weights of the other Centres allows the federated model to remain robust in front of optimization errors during training.
@@
Furthermore, in order to support the validity of the proposed training strategy, the results obtained through alternative well-established training methodologies using the Li et al. model were presented and discussed in detail in B, demonstrating the effectiveness of the chosen Random Multi-2D training approach.
This work established the foundations for future developments. Several limitations of the employed architecture could be addressed by adopting more advanced medical image synthesis methodologies [5, 4, 28], optimising their efficiency, and ensuring compatibility with the FL environment. A possible improvement could also involve optimizing the implementation of 3D architectures for federated scenarios by leveraging more efficient training strategies, such as model pruning, knowledge distillation, or dimensionality reduction [56]. Additionally, exploring hybrid federated approaches, such as cross-silo horizontal FL with asynchronous updates or client clustering [47, 67], could reduce synchronization times.
A further limitation of the present study was the available computational resources, which imposed a limit on the number of clients that could be incorporated into the federation. In future work, a real-world deployment could be considered to mitigate the constraints imposed by a single, although powerful, machine. In addition, the evaluation of the generated sCTs using the emerging sCT quality control algorithms [72] has the potential to enhance the comprehension of the limitations of the federated framework and to introduce enhancements in future studies.
Moreover, the implementation of specific pre-processing or data harmonisation techniques can be considered to enhance the homogeneity of the data and to mitigate potential hidden artefacts [20]. This, consequently, will improve the quality of the sCTs for both the Centres within the federation and in the generalisation task.
@@ Acknowledgments
Appendix A Investigation of deep learning architectures
This preliminary investigation was conducted using data from centres A and C, with Centre E used as the unseen dataset. This choice reduced computational time and allowed evaluation of multiple architectures. The hyper-parameters used are detailed in Section 4.2. The minimal augmentation pipeline was used to further reduce the computational time. Among the selected architectures, were: U-Net (described in Section 3.3), Conditional GAN (Pix2Pix GAN, with a U-Net-based generator) [26], and CycleGAN (with a ResNet-based generator) [73], which are commonly used in image-to-image translation or CT synthesis tasks [13, 57, 8]. Despite its inherent complexity, we also evaluated the winning deep learning (DL) model of the SynthRAD challenge, which used a patch-based approach for centralized MRI-to-sCT [25], namely the MSEP architecture [11].
Pix2Pix Conditional GAN. The Pix2Pix conditional GAN (cGAN) model employed a U-Net-based generator and a PatchGAN discriminator, used to have a $30\times 30$ probability patch in output rather than a single value. For this model, an hybrid loss that integrated the pixel-wise error of the prediction and the discriminator’s assessment was employed.
CycleGAN. The CycleGAN used in this study was composed of two generators and two discriminators. The generator networks were implemented as instances of a ResNet, while PatchGANs with an output size of $30\times 30$ were employed for the discriminators. The generator loss was calculated as a weighted average of three components: paired pixel-wise loss, cycle-consistency loss, and adversarial loss.
MSEP. To enable its use in a FL context, we adapted the original implementation accordingly. To train the MSEP, patches with dimensions $48\times 160\times 160$ were extracted from each sample. The location of each patch was chosen randomly every epoch to increase the amount of possible training input data needed by intrinsic nature of the architecture. Since with this 3D DL approach, the model predicted overlapping patches and the number of predictions was not constant across the volume, the average value was calculated for each voxel using the median voting approach to reach a consensus.
In this scenario, the hardware needed to lead the experiment was an NVIDIA A100 80GB GPU, 32 CPU cores, and 256GB of RAM from a Supermicro AS-4124GS-TNR 4U Rackmount GPU SuperServer.
Center E (Unseen)
Model MAE [HU] SSIM PSNR [dB] Time
U-Net 110.4 (91.2-126.1) 0.88 (0.85-0.89) 26.44 (25.42-27.27) 6h 00min
Pix2Pix 122.0 (104.6-139.1) 0.87 (0.85-0.87) 26.31 (25.27-26.98) 4h 00min
CycleGAN 121.4 (115.7-130.9) 0.87 (0.85-0.88) 26.30 (25.32-27.07) 10h 28min
MSEP 240.5 (220.7-250.7) 0.79 (0.77-0.80) 23.70 (23.09-24.29) 85h 37min*
Table 4: The median and interquartile range (IR) for the mean absolute error (MAE), the structural similarity index (SSIM), and the peak signal-to-noise ratio (PSNR) obtained using different architectures, as well as the total time required for the experiments.
(*) Time required to perform a single round.
The results reported in 4 demonstrated that the U-Net obtained the optimal overall performance, taking into account both image similarity metrics and computational efficiency, with a training duration of 6 hours employing only Center A and C.
Although the timing of Pix2Pix GAN was favourable, its conversion performance was inferior to that of U-Net under the same conditions. This difference was especially evident from the median and interquartile range of the MAE presented in Table 4.
The CycleGAN model demonstrated inferior performance in terms of image similarity when employing equivalent settings, with a training duration of 10 hours and 28 minutes that was nevertheless less efficient than the U-Net.
Regarding the MSEP transformer-based architecture, despite efforts to reduce computational overhead — such as limiting the number of participating clients and reducing the size of the 3D patches — the time required for a single federated round remained prohibitively high. Specifically, one round took approximately 85 hours and 37 minutes. Moreover, the federated MSEP architecture’s performance fell short when compared to its centralised counterpart. This decline can be primarily attributed to: (i) the adjustments made to the training pipeline to accommodate the federated framework, and (ii) the reduced number of rounds and clients necessitated by the computational complexity of the 3D approach.
Appendix B Training methods investigation
The primary objective of this investigation was to further evaluate the effectiveness of the proposed FL training approach compared to well-established training methods described in the literature for the MRI-to-sCT task in a centralised setting, described in Section 2.1. This investigation was applied to the selected model, U-Net by Li et al. [41], which had previously demonstrated better performance in terms of efficiency when compared with different U-Nets. The hyper-parameters used are detailed in Section 4.2. The model was trained across Centres A, B, C, D, with Centre E serving as an unseen test dataset. The performance of each training methodology was assessed using the metrics described in Section 4.3, computed on the unseen dataset. In the patch-based method, we extracted randomly cropped patches of $160\times 160$ pixels to maximise contextual information. Overlapping regions were merged using pixel-wise averaging.
Center E (Unseen)
Method MAE [HU] SSIM PSNR [dB] Time
2D+ 104.0 (99.5-115.6) 0.89 (0.86-0.89) 26.49 (25.43-27.35) 16h 00min
2D Patches 162.4 (155.3-173.0) 0.84 (0.82-0.85) 25.52 (24.65-26.12) 8h 39min
Multi-2D 111.7 (107.7-122.3) 0.88 (0.85-0.89) 26.33 (25.32-27.20) 9h 35min
Random
Multi-2D (Our) 102.0 (96.7-110.5) 0.89 (0.86-0.89) 26.58 (25.52-27.42) 12h 34min
Table 5: The median and interquartile range (IR) for the mean absolute error (MAE), structural similarity index (SSIM), and peak signal-to-noise ratio (PSNR) obtained for each training paradigm. The total training time required to complete 25 rounds is also reported.
The results presented in Table 5 further supported the superiority of the proposed Random Multi-2D approach over the alternative 2D methods evaluated. Specifically, the Random Multi-2D method achieved performance in terms of image similarity metrics comparable to the 2D+ approach, where each model was specifically trained on a single plane (axial, coronal and sagittal). Our training method required only a single model for all anatomical planes, thereby simplifying the federated training and inference processes. Additionally, the training time was significantly reduced with the Random Multi-2D model completing training in 12 hours and 34 minutes, compared to 16 hours for the 2D+ models. In contrast, the 2D patch-based method requiring less training time but exhibited inferior performance across all image similarity metrics.
In comparison with the classical Multi-2D [57] method, the proposed Random Multi-2D approach yielded superior image similarity metrics, albeit at a higher computational cost, which can be attributed to the shuffling operation conducted by local clients during federated training. Moreover, the utilisation of our method enabled the federated model to attain superior outcomes in comparison to alternative training methods after only 5 rounds.
The findings of this additional study demonstrated that the Random Multi-2D paradigm not only exhibited competitive performance in comparison to the 2D+ approach, but also demonstrated improved computational efficiency and substituting a more generic model for three specific models. Furthermore, the proposed training method yielded superior generalisation performance compared to the classical Multi-2D method, thus justifying the suitability of the Random Multi-2D paradigm for FL MRI-to-sCT scenarios.
Appendix C Centre B artefact measurement
@@
Appendix D Aggregation strategies investigation
Convergence trends
@@ Appendix D Aggregation strategies investigation
Generated sCTs
As shown in Table 3, the evaluated strategies produced qualitatively comparable results, with similar mean absolute errors (MAE). This observation is further supported by Figure 11, which provides a visual comparison of the results obtained using different aggregation strategies on an axial slice of an example sCT from the same patient of the unseen Centre E. The discrepancy between the ground truth CT and the sCT is also illustrated. Notably, the sCT corresponding to FedYogi demonstrated a more pronounced discrepancy from the ground truth CT, in line with the highest MAE value reported in Table 3. This finding confirmed the limitations of FedYogi for the MRI-to-sCT task compared to the other aggregation strategies, as well as the pivotal role of the aggregation strategy in the federated context. Indeed, as well as determining the quality of the generated sCT, the strategy had a decisive influence on model convergence.
Figure 11: Visual comparison of sCT generated with different aggregation strategies on an axial slice from a patient of Centre E. The discrepancy between the ground truth CT and the sCT is highlighted on the right.